OpenStudy (anonymous):

3.986699933448 Correct to six decimal places? Round

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} 3.98669&{\color{red}{ 9}}933448\\ &\uparrow \\ &6^{th}\quad decimal \end{array}\)

OpenStudy (anonymous):

I understand what the sixth decimal is. I need this rounded. When I stick that answer in my onine hw, it says it is incorrect.

OpenStudy (whpalmer4):

You need to find the 6th place so that you know where the 7th place is. It's the value of the 7th place that determines the answer — if it is 0-4, you just copy what you have through the 6th place. If it is 5-9, you copy what you have through the 6th place, then adjust the 6th place up by 1 (which may adjust the other places as well).

OpenStudy (jdoe0001):

ahemm I see... right... well... \(\large \begin{array}{llll} 3.986&{\color{red}{699}} 33448\\ 3.986& {\color{red}{ 700}}33448\\ 3.986&70000000\\ 3.986700 \end{array}\)

OpenStudy (jdoe0001):

hmm ohh.. .. yes :/

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} 3.986&{\color{red}{699}}933448\\ 3.986& {\color{red}{ 700}}933448\\ 3.986&70000000\\ 3.986700 \end{array}\)

OpenStudy (anonymous):

Thanks!